### v2

@@ 6 Conclusions
In this paper, we have considered the scenario where a flying AU is paired with a static TU through power-domain uplink NOMA. Despite many existing works, we have investigated the performance of finite-sample channel estimation procedures. We have shown that blind and pilot-aided techniques can be suitably combined to accurately estimate the AU and TU channel that are characterized by different second-order wide-sense properties in both the ACS (or Doppler) and modulation domains. Specifically, the AU mobility induces ACS features in the signal received from sky due to Doppler effects that are not present in the signal received from ground, provided that the TU is static. Moreover, the AU can employ a noncircular modulation format, as opposed to the TU that transmits circular symbols. We have demonstrated that Doppler shifts and time delays of the AU can be estimated blindly with high accuracy that is basically independent of the TU and noise powers. The remaining AU parameters, i.e., channel gains and AoAs, can be instead estimated through a pilot-based LS approach that exploit the previously acquired knowledge of the Doppler shifts and time delays of the AU. By equipping the BS with the CSI knowledge of the AU, the overall channel of the TU can be subsequently acquired through a high-precision pilot-based BWLU estimator.

### v4

@@ 6 Conclusions and directions for future work
In this paper, we have considered the scenario where a flying AU is paired with a static TU through power-domain uplink NOMA. Despite many existing works, we have investigated the performance of finite-sample channel estimation procedures. We have shown that blind and pilot-aided techniques can be suitably combined to accurately estimate the AU and TU channel that are characterized by different second-order wide-sense properties in both the ACS (or Doppler) and modulation domains. Specifically, the AU mobility induces ACS features in the signal received from sky due to Doppler effects that are not present in the signal received from ground, provided that the TU is static. Moreover, the AU can employ a noncircular modulation format, as opposed to the TU that transmits circular symbols. We have demonstrated that Doppler shifts and time delays of the AU can be estimated blindly with high accuracy that is basically independent of the TU and noise powers. The remaining AU parameters, i.e., channel gains and AoAs, can be instead estimated through a pilot-assisted LS approach that exploits the previously acquired knowledge of the Doppler shifts and time delays of the AU. By equipping the BS with the CSI knowledge of the AU, the overall channel of the TU can be subsequently acquired through a high-precision pilot-based BWLU estimator.
The performance of the proposed estimators has been corroborated through Monte Carlo numerical results. In this respect, an interesting research subject consists of theoretically investigating the consistency and asymptotic distribution of such algorithms. Another appealing research issue is to extend our framework to the case in which TUs also communicate with the BS according to a NOMA scheme.